a metal solid at fahrenheit below 1772.7 determine (in terms of an inquality) those celsius temperature for which the metal stays a solid. Use formula F=9/5 c + 32
Answer by stanbon(57222)

a metal solid at fahrenheit below 1772.7 determine (in terms of an inquality) those celsius temperature for which the metal stays a solid.
Use formula F=9/5 c + 32
------
1772.7 = (9/5)c + 32
-------
(9/5)c = 1740.7
---
c = 967.056
================
Cheers,
Stan H.
================